OpenStudy (anonymous):

given the midpoint and one endpoint of a line segment, find the other endpoint

OpenStudy (anonymous):

the distance between one endpoint and midpoint is same as distance to other ends point(from midpoint)

OpenStudy (anonymous):

how do you do it

OpenStudy (anonymous):

\[(\frac{x_1 +x_2}{2},\frac{y_1+y_2}{2})=(x,y)\]

OpenStudy (anonymous):

^^^^Use this.

OpenStudy (anonymous):

formula for mid point

OpenStudy (anonymous):

how do you do it with the endpoint

OpenStudy (anonymous):

reverse way?

OpenStudy (anonymous):

just plug in the values u have and solve for the 1 u dont

OpenStudy (anonymous):

x,y are midpoints \(x_1,y_1\:\:\:\:\: and \:\:\:x_2,y_2\) are end-points.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

midpoint x value: \[x_{midpoint}=(1/2)(x_{endpoint1}+x_{endpoint2})\]

OpenStudy (anonymous):

thx guys

OpenStudy (anonymous):

midpoint (c,d) end points (x1, y1) (x2,y2) \[\left(x_1-c\right),\left(y_1-d\right)\] \[x_2=c+\left(\left(x_1-c\right)\right)\] \[y_2=d+\left(\left(y_1-d\right)\right)\]

OpenStudy (across):

We three points: A, B, and C, where A and B are endpoints and C is a midpoint. Then: \[B=(x_{B}, y_{B})=(2x_{C}-x_{A}, 2y_{C}-y_{A}).\]